OpenStudy (anonymous):

using elimination or substitution solve this: 2y^2+ 3x=33 and x+4y+7=0

OpenStudy (anonymous):

Substitute " X=-(4Y+7) " into the other one

OpenStudy (anonymous):

you'll get :2y^2-3(4y+7)=33 then 2y^2-12y-21-33=0 y^2-6y-27=0 (y-9)(y+3)=0 y=9or-3

OpenStudy (anonymous):

x=-(4y+7) so x=-(4x9+7) or x=-(4x(-3)+7)

OpenStudy (anonymous):

ans: x=-43,y=9 or x=5,y=-3

OpenStudy (anonymous):

Thank you so much :)